OpenStudy (anonymous):

If the sequence is geometric, give its common ratio. 7, 49, 343, 2401 ... 42 9 7 not geometric

OpenStudy (anonymous):

2401/ 343 = 7 343/49 = 7 48/7 = 7 Notice a pattern. this is essentially what you do to find the common ratio.